### v2

# Strong nebular emissions associated with Mg ii absorptions detected in the SDSS spectra of background quasars

Abstract
We present long-slit spectroscopic observations of 40 Galaxy On Top Of Quasars (GOTOQs) at using the South African Large Telescope. Using this and available photometric data, we measure the impact parameters of the foreground galaxies to be in the range of 316 kpc with a median value of 8.6 kpc. This is the largest sample of galaxies producing Mg ii absorption at such low impact parameters. These quasar-galaxy pairs are ideal for probing the disk-halo interface. At such impact parameters, we do not find any anti-correlation between rest equivalent width (REW) of Ca ii, Mn ii, Fe ii, Mg ii and Mg i absorption and impact parameter. These sight lines are typically redder than those of strong Mg ii absorbers, with the color excess, E(BV) for our sample ranging from 0.191 to 0.422, with a median value of 0.058. In the E(BV) vs. W3935 plane, GOTOQs occupy the same region as Ca ii absorbers. For a given E(BV), we find larger W3935 than what has been found in the Milky Way, probably due to a smaller dust-to-gas ratio in GOTOQs. Galaxy parameters could be measured for twelve cases, and their properties seem to follow the trends found for strong Mg ii absorbers. Measuring the host galaxy properties for the full sample using HST photometry or AO-assisted ground-based imaging is important to gain insights into the relationship between the stellar mass of galaxies and the metal line REW distributions at low impact parameters.
keywords:
galaxies: evolution – galaxies: groups: general – galaxies: haloes – galaxies: high-redshift – quasars: absorption lines.1 Introduction
Absorption lines imprinted in the spectra of bright background quasars are very useful in probing the physical conditions and the kinematics of the gas in the galactic disk and/or the circumgalactic medium (CGM) of foreground galaxies. In particular, absorption produced by the Mg ii doublet has been used to identify intervening galaxies (Bergeron & Boissé, 1991; Steidel, 1995). Systematic studies over the past couple of decades and recent advances in the IFU-based spectroscopy (see for example, Bouché et al., 2012; Nielsen et al., 2013; Schroetter et al., 2019; Dutta et al., 2020; Huang et al., 2021; Lundgren et al., 2012, 2021), have allowed us to establish the relationship between Mg ii absorbers and host galaxies (or group of galaxies) over a wide range of impact parameters (D). However, these studies mostly probe large impact parameters, and we have very little statistics on the properties of Mg ii absorbers at, say, kpc (see Kacprzak et al., 2013, however). For example, out of about 560 unique quasar-galaxy pairs studied in various big surveys of Mg ii absorption like MAGIICAT (Nielsen et al., 2013), M3 halo project (Huang et al., 2021) and MAGG (Dutta et al., 2020), only 5 have an impact parameter of less than ten kpc. Kacprzak et al. (2012) have reported seven Mg ii systems with kpc, but most of these quasar-galaxy paris are restricted to very low redshifts (i.e., ).
A quasar line of sight passing very close to the center of an intervening galaxy (say with ) is most likely to probe the disk or disk-halo (or CGM) interface of the galaxy. Over the past few years, there have been significant explorations of the disk-halo interface of our own Galaxy (Bish et al., 2019; Hobbs & Feldmann, 2020) as well as some of the low-redshift galaxies (Boettcher et al., 2019). These studies confirm the presence of inflows and outflows along the disk-halo interface consistent with the ‘galactic fountain’ models (Shapiro & Benjamin, 1991; Rubin et al., 2022). Probing gas flows at such scales is important for understanding the sustained star formation in galaxies.
The best strategy for identifying high- quasar-galaxy pairs with low impact parameters is to look for the nebular emission lines from the intervening galaxy in the spectrum of the background quasar. If a star-forming foreground galaxy is present within an angular separation of (for example, in the case of SDSS DR-12) or (for example in the case of SDSS DR-7) to the background quasar, the fiber used to observe the background quasar will pick the nebular line emissions from the foreground galaxy. This is now proven to be a very efficient technique for identifying host galaxies of Mg ii absorption systems situated at very small angular separations (Noterdaeme et al., 2010; York et al., 2012; Straka et al., 2015; Joshi et al., 2018; Kulkarni et al., 2022). Such systems are referred to as Galaxies On Top of Quasars (or GOTOQs in short - York et al., 2012). Starting from the Mg ii absorption doublet present on the background quasar spectra, Joshi et al. (2017) have identified 198 GOTOQs over the redshift range . However, note that measuring the impact parameters of the foreground galaxies directly from the fiber-fed spectroscopy is not possible.
In Guha & Srianand (2023), we have investigated the nature of these GOTOQs using multiband images available from the Dark Energy Spectroscopic Instrument (DESI) Legecy Survey(Dey et al., 2015, 2019). There are 74 GOTOQs in the sample of Joshi et al. (2017) for which photometric extension around the quasar image was detected in deep images, after subtracting the quasar contribution Guha & Srianand (2023) were able to measure the impact parameters of the galaxies to be in the range . This has substantially increased the number of Mg ii absorbing host galaxies probed at such low impact parameters and allowed us to measure of Å and an exponential scale length of 21.6 kpc for the relationship. While the photometric redshift estimated for the residual light (after subtracting the quasar contribution) is consistent (within ) with the redshift of the Mg ii absorbers, independent confirmation using the spatially resolved galaxy spectra is unavailable.
One of the primary objectives of this study is to verify that the photometric extensions observed in GOTOQs correspond to the host galaxies of Mg ii absorption. Another aim is to identify the locations of galaxies in cases where there are no clear extensions visible in the available photometric images. We use long-slit spectroscopy using the Southern African Large Telescope (SALT, Buckley et al., 2005) with slit orientated along different position angles for this purpose.
This paper is organized as follows. Section 2 describes the sample constructed to study the host galaxies of strong Mg ii absorption systems at small impact parameters, and section 3 describes settings used for our SALT observations and data reduction. Section 4 describes our main results, while section 5 summarizes the discussions and conclusions of this paper. Throughout this paper, we use a flat cosmology with Hubble constant, = 70 and = 0.3.
2 The GOTOQS sub-sample
(1) | (2) | (3) | (4) | (5) | (6) | (7) | (8) | (9) | (10) | (11) |
No. | Quasar | Plate | MJD | Fiber | E(BV | |||||
1 | BOSS-J000123.61032317.9 | 4354 | 55810 | 0282 | 3.1210 | 0.8918 | 0.8922 | 33.67 4.75 | 2.12 0.28 | 0.0620.004 |
2 | BOSS-J005436.35+005055.6 | 4224 | 55481 | 0746 | 2.2363 | 0.4810 | 0.4812 | 14.36 0.74 | 1.87 0.37 | 0.0900.003 |
SDSS-J005436.35+005055.6 | 0692 | 52201 | 0539 | 2.2386 | 0.4806 | 0.4813 | 15.55 2.56 | 1.91 0.75 | 0.0840.002 | |
3 | SDSS-J005408.46094638.2 | 0657 | 52177 | 0515 | 2.1232 | 0.4778 | 0.4778 | 15.81 2.30 | 2.41 0.15 | 0.0760.001 |
4 | BOSS-J012916.14003211.4 | 4229 | 55501 | 0182 | 1.8370 | 0.8991 | 0.8997 | 31.44 3.30 | 3.50 0.26 | 0.0480.002 |
5 | BOSS-J013132.92040137.5 | 7048 | 56575 | 0710 | 2.2700 | 0.5017 | 0.5017 | 17.78 2.56 | 1.70 0.12 | 0.0850.002 |
6 | SDSS-J020317.22010122.8 | 0701 | 52179 | 0087 | 1.7210 | 0.7253 | 0.7258 | 26.10 7.53 | 2.10 0.51 | 0.0500.005 |
7 | BOSS-J023529.82+005916.3 | 3744 | 55209 | 0952 | 2.7160 | 0.9250 | 0.9245 | 34.07 5.42 | 0.93 0.43 | 0.0510.006 |
8 | BOSS-J023608.18+005604.0 | 6781 | 56274 | 0791 | 0.9680 | 0.8212 | 0.8216 | 19.70 2.55 | 2.02 0.12 | 0.0110.002 |
SDSS-J023608.18+005604.0 | 0408 | 51821 | 0321 | 0.9671 | 0.8212 | 0.8216 | 10.67 3.56 | 1.77 0.27 | 0.0370.003 | |
9 | SDSS-J023618.98000529.1 | 0408 | 51821 | 0237 | 0.9806 | 0.6448 | 0.6460 | 17.16 4.86 | 1.26 0.40 | 0.0790.004 |
10 | SDSS-J030730.61074555.6 | 0459 | 51924 | 0309 | 0.7543 | 0.6986 | 0.6994 | 28.81 4.00 | 2.34 0.27 | 0.1000.002 |
11 | BOSS-J080751.03+094554.7 | 4510 | 55559 | 0081 | 2.5900 | 0.7470 | 0.7470 | 24.17 2.79 | 2.12 0.21 | 0.0350.003 |
12 | SDSS-J080808.67+064108.7 | 1756 | 53080 | 0184 | 2.1133 | 0.4326 | 0.4326 | 8.96 2.31 | 1.96 0.59 | 0.0070.003 |
13 | SDSS-J084915.67+065013.0 | 1189 | 52668 | 0322 | 2.1970 | 0.8432 | 0.8437 | 140.24 4.25 | 2.60 0.14 | 0.0560.004 |
BOSS-J084915.67+065013.0 | 5289 | 55893 | 0128 | 2.1970 | 0.8432 | 0.8438 | 125.55 6.90 | 3.00 0.27 | 0.0420.004 | |
14 | SDSS-J085051.97+083026.6 | 1299 | 52972 | 0333 | 0.6635 | 0.3767 | 0.3767 | 5.56 1.74 | 1.75 0.15 | 0.0540.003 |
15 | SDSS-J085830.53+055606.0 | 1191 | 52674 | 0283 | 1.8524 | 0.4318 | 0.4318 | 17.76 1.62 | 2.90 1.03 | 0.0520.004 |
16 | SDSS-J090830.32+042436.3 | 1192 | 52649 | 0019 | 1.7518 | 0.4362 | 0.4363 | 5.02 2.01 | 1.69 0.37 | 0.0010.002 |
17 | BOSS-J091709.35+033734.0 | 3820 | 55542 | 0008 | 2.1060 | 0.6622 | 0.6620 | 18.82 4.78 | 2.66 0.80 | 0.0300.003 |
18 | BOSS-J092915.02-012621.0 | 3781 | 55243 | 0978 | 2.2500 | 0.6990 | 0.7002 | 8.57 1.85 | 1.47 0.43 | 0.1910.004 |
19 | SDSS-J094222.51+001641.9 | 0266 | 51630 | 0396 | 1.4065 | 0.5243 | 0.5247 | 20.86 2.80 | 2.44 0.23 | 0.1000.002 |
20 | BOSS-J101446.63+083500.9 | 5334 | 55928 | 0860 | 2.5120 | 1.0017 | 1.0012 | 49.27 8.38 | 4.75 0.08 | 0.0730.002 |
SDSS-J101446.64+083500.9 | 3155 | 54829 | 0179 | 2.5113 | 1.0017 | 1.0011 | 46.53 8.43 | 4.39 0.10 | 0.0930.003 | |
21 | BOSS-J102518.88001755.6 | 3832 | 55289 | 0388 | 2.2850 | 0.8801 | 0.8803 | 39.53 6.84 | 3.82 0.29 | 0.1070.002 |
22 | BOSS-J111755.71+075932.2 | 5369 | 56272 | 0393 | 2.6350 | 0.7681 | 0.7687 | 24.01 2.72 | 1.98 0.61 | 0.0120.005 |
23 | SDSS-J120908.53+022734.3 | 0517 | 52024 | 0541 | 1.3370 | 0.6695 | 0.6695 | 59.08 3.97 | 3.27 0.22 | 0.0330.003 |
24 | SDSS-J121242.26+083410.9 | 1624 | 53386 | 0453 | 1.5368 | 0.5330 | 0.5322 | 12.56 4.88 | 2.12 0.20 | 0.0180.002 |
25 | BOSS-J121838.10+030900.1 | 4750 | 55630 | 0774 | 2.2510 | 0.5222 | 0.5220 | 11.53 1.85 | 2.80 0.39 | 0.0140.002 |
26 | BOSS-J123514.36+030416.7 | 4754 | 55649 | 0592 | 2.1950 | 0.5313 | 0.5338 | 17.82 2.37 | 1.36 0.23 | 0.1250.002 |
27 | BOSS-J131454.20+033042.3 | 4760 | 55656 | 0096 | 2.7000 | 0.7792 | 0.7801 | 18.72 2.71 | 1.84 0.21 | 0.4220.003 |
28 | BOSS-J131918.88030621.2 | 4049 | 55591 | 0418 | 2.3400 | 0.6772 | 0.6777 | 32.77 6.30 | 2.06 0.47 | 0.0600.003 |
29 | SDSS-J135445.72+075637.6 | 1806 | 53559 | 0043 | 1.3235 | 0.7126 | 0.7119 | 26.39 3.95 | 2.15 0.31 | 0.0950.003 |
30 | SDSS-J135634.16+060133.6 | 1805 | 53875 | 0013 | 1.3929 | 0.6292 | 0.6298 | 29.37 10.55 | 2.52 0.19 | 0.0070.002 |
31 | SDSS-J140446.28+062340.2 | 1824 | 53491 | 0231 | 1.1852 | 0.4515 | 0.4521 | 16.11 1.85 | 2.28 0.33 | 0.1420.002 |
32 | BOSS-J142232.38003043.9 | 4028 | 55621 | 0422 | 2.7060 | 1.0127 | 1.0127 | 40.50 13.70 | 1.59 0.13 | 0.0140.003 |
33 | SDSS-J142953.20+041551.6 | 0585 | 52027 | 0473 | 1.9410 | 0.7134 | 0.7126 | 22.40 4.21 | 3.57 0.33 | 0.1160.003 |
34 | SDSS-J161728.95+061604.3 | 1731 | 53884 | 0113 | 1.2501 | 0.7884 | 0.7878 | 90.74 5.72 | 3.20 0.18 | 0.0970.002 |
35 | BOSS-J215059.09+022557.2 | 5146 | 55831 | 0878 | 2.3780 | 0.5682 | 0.5673 | 7.01 1.45 | 2.37 0.92 | 0.0520.003 |
36 | BOSS-J223244.93+011058.1 | 4203 | 55447 | 0804 | 1.0090 | 0.4899 | 0.4901 | 4.20 0.96 | 1.69 0.37 | 0.1850.002 |
37 | BOSS-J224405.09021547.9 | 4363 | 55537 | 0778 | 2.5740 | 0.9243 | 0.9246 | 18.73 2.92 | 3.77 0.20 | 0.0880.002 |
38 | BOSS-J225028.85+031146.4 | 4293 | 55509 | 0453 | 2.5340 | 0.7198 | 0.7196 | 17.31 2.97 | 2.67 0.12 | 0.0120.002 |
39 | SDSS-J234257.50+010121.3 | 1903 | 53357 | 0420 | 2.4520 | 0.6767 | 0.6764 | 11.24 2.78 | 3.10 0.38 | 0.0570.004 |
BOSS-J234257.50+010121.3 | 4213 | 55449 | 0912 | 2.4520 | 0.6768 | 0.6763 | 8.45 3.23 | 2.35 0.27 | 0.0740.003 | |
40 | BOSS-J235639.31040614.4 | 7034 | 56564 | 0506 | 2.8800 | 0.7709 | 0.7705 | 55.89 8.51 | 3.63 0.34 | 0.0620.005 |
In this work, we mainly focus on a sub-sample of GOTOQs from the original sample of Joshi et al. (2017). In brief, Joshi et al. (2017) inspected all the quasar spectra with intervening Mg ii absorption listed in the SDSS Fe ii /Mg ii metal absorber catalog (Zhu & Ménard, 2013) and searched for associated nebular emissions on top of the quasar spectra, utilizing either the [O ii] line with a minimum detection significance of or each of the [O iii] 4960, 5008 lines with significances. This resulted in 198 GOTOQs in the redshift interval 0.35 1.1. As our primary focus of this study is to precisely calculate the impact parameter of the foreground Mg ii absorption host galaxies using the long-slit spectra obtained at different position angles (PAs), while constructing the GOTOQs sub-sample for this work, we only considered systems that are accessible to the SALT, i.e., the declination of the systems being less than .
We introduced a cut-off on the [O ii] luminosity to limit the number of sources. We only consider those systems for which the luminosity [O ii] emission line detected in SDSS quasar spectra are at least = . Given the SFR calibrations (Kennicutt, 1998), this threshold in [O ii] line luminosity corresponds to a minimum SFR (uncorrected for dust extinction and fiber loss) of . However, note that the actual SFR can be much higher since we did not correct the fiber effects and the extinctions due to interstellar dust. The two criteria discussed above resulted in 53 GOTOQs, out of which we completed long-slit observations for 40 systems using SALT. The details of the GOTOQs observed with SALT are presented in Table 1. The Mg ii rest equivalent width, of the observed sample varies between 0.93Å to 4.75Å with a median value of 2.21Å.
To compare this GOTOQs sub-sample against the original sample, in Figure 1, we plot the histograms of both these systems based on various absorption and emission line properties. The left panel shows the histograms of [O ii] line luminosity for our GOTOQs sub-sample (blue) and the entire sample (orange). By construction, this sub-sample should correspond to the more luminous [O ii] systems, as evident from this plot. For the original sample [O ii] luminosity ranges with a median value of while for the sub-sample the [O ii] luminosity ranges with a median value of .
The middle panel shows the histograms of for our GOTOQs sub-sample (blue) and the entire sample (orange). Compared to the range of with a median value of 2.38Å for the full sample, for our sample ranges with a median value of 2.21Å. The plot shows that the distributions of for the two samples are similar. The KS test yields that these two distributions are drawn from the same underlying population ( with a value of 0.85).
Finally, the right panel shows the histograms of the Mg ii absorption redshift for our GOTOQs sub-sample (blue) and the complete sample (orange). The absorption redshift for our sample and the complete sample varies over the range 0.37 1.01 with a median value of 0.69 and 0.36 1.1 with a median value of 0.66 respectively. KS test yields that the redshift distributions of these two samples are likely drawn from the same distributions ( with a value of 0.90). In summary, the GOTOQ sub-sample used in this work probes the GOTOQs systems over the complete and redshift ranges with slightly higher compared to the primary GOTOQs sample.
3 Observations and data reductions
No. | Quasar | Date | Exposure (s) | PA (deg.) | Grating Angel (deg.) | Wavelength Range (Å) |
(1) | (2) | (3) | (4) | (5) | (6) | (7) |
1 | J000123.61032317.9 | 2017-06-26 | 2480 | 90 | 21.5 | 6560 – 9510 |
2017-07-01 | 2480 | 90 | 21.5 | 6560 – 9510 | ||
2023-07-17 | 2480 | 150 | 21.125 | 6420 – 9370 | ||
2 | J005408.46094638.2 | 2018-08-12 | 2500 | 45 | 17.375 | 5050 – 8050 |
2018-10-19 | 2600 | 315 | 17.375 | 5050 – 8050 | ||
3 | J005436.35+005055.6 | 2023-07-17 | 2500 | 75 | 17.0 | 4900 – 7900 |
4 | J012916.14003211.4 | 2017-07-27 | 2480 | 315 | 21.5 | 6560 – 9510 |
2017-08-28 | 2480 | 315 | 21.5 | 6560 – 9510 | ||
5 | J013132.92040137.5 | 2023-08-04 | 2500 | 15 | 17.375 | 5050 – 8050 |
6 | J020317.22010122.8 | 2017-08-03 | 2500 | 135 | 19.625 | 5860 – 8860 |
2017-09-13 | 2480 | 45 | 19.625 | 5860 – 8860 | ||
7 | J023529.82+005916.3 | 2017-07-29 | 2520 | 60 | 21.5 | 6560 – 9510 |
2017-10-17 | 2400 | 330 | 21.5 | 6560 – 9510 | ||
8 | J023608.18+005604.0 | 2023-09-19 | 2480 | 90 | 20.75 | 6280 – 9250 |
2023-09-12 | 2480 | 0 | 20.75 | 6280 – 9250 | ||
9 | J023618.98000529.1 | 2023-10-16 | 2500 | 100 | 18.875 | 5590 – 8590 |
10 | J030730.61074555.6 | 2022-10-27 | 2400 | 0 | 19.625 | 5860 – 8860 |
2022-10-01 | 2400 | 90 | 19.625 | 5860 – 8860 | ||
11 | J080751.03+094554.7 | 2023-12-19 | 2200 | 80 | 20.0 | 6005 – 8985 |
2023-12-22 | 2200 | 170 | 20.0 | 6005 – 8985 | ||
12 | J080808.67+064108.7 | 2023-12-16 | 2300 | 90 | 16.625 | 4780 – 7780 |
13 | J084915.67+065013.0 | 2022-02-23 | 2300 | 63 | 21.125 | 6420 – 9370 |
2022-02-26 | 2300 | 162 | 21.125 | 6420 – 9370 | ||
14 | J085051.97+083026.6 | 2024-04-03 | 2240 | 45 | 16.25 | 4620 – 7650 |
15 | J085830.53+055606.0 | 2022-12-18 | 2560 | 112 | 16.625 | 4780 – 7780 |
2023-01-23 | 2400 | 22 | 16.625 | 4780 – 7780 | ||
16 | J090830.32+042436.3 | 2024-04-12 | 2240 | 135 | 16.625 | 4780 – 7780 |
17 | J091709.35+033734.0 | 2024-04-04 | 2500 | 90 | 18.875 | 5590 – 8590 |
18 | J092915.02012621.0 | 2023-12-13 | 2500 | 90 | 19.25 | 5725 – 8725 |
19 | J094222.51+001641.9 | 2024-03-06 | 2500 | 90 | 17.75 | 5170 – 8190 |
20 | J101446.63+083500.9 | 2022-02-06 | 2240 | 177 | 19.625 | 5860 – 8860 |
21 | J102518.88001755.6 | 2022-12-30 | 2560 | 0 | 21.5 | 6560 – 9510 |
2022-04-20 | 2540 | 77 | 21.5 | 6560 – 9510 | ||
22 | J111755.71+075932.2 | 2024-03-05 | 2240 | 315 | 20.0 | 6005 – 8985 |
23 | J120908.53+022734.3 | 2018-02-16 | 2400 | 0 | 19.25 | 5725 – 8725 |
2018-02-22 | 2400 | 90 | 19.25 | 5725 – 8725 | ||
24 | J121242.26+083410.9 | 2024-04-13 | 2200 | 0 | 17.375 | 5050 – 8050 |
25 | J121838.10+030900.1 | 2024-02-12 | 2500 | 315 | 17.75 | 5165 – 8195 |
26 | J123514.36+030416.7 | 2023-04-25 | 2560 | 0 | 17.75 | 5165 – 8195 |
27 | J131454.20+033042.3 | 2024-03-09 | 2500 | 90 | 20.375 | 6150 – 9120 |
2024-03-10 | 2100 | 0 | 20.375 | 6150 – 9120 | ||
28 | J131918.88030621.2 | 2017-06-16 | 2500 | 45 | 19.25 | 5725 – 8725 |
2017-07-16 | 2335 | 315 | 19.25 | 5725 – 8725 | ||
29 | J135445.72+075637.6 | 2017-05-18 | 2200 | 315 | 19.625 | 5860 – 8860 |
2017-05-18 | 2200 | 45 | 19.625 | 5860 – 8860 | ||
30 | J135634.16+060133.6 | 2017-04-28 | 2200 | 277 | 18.875 | 5590 – 8590 |
2017-05-20 | 2200 | 20 | 18.875 | 5590 – 8590 | ||
31 | J140446.28+062340.2 | 2023-07-17 | 2300 | 0 | 16.625 | 4780 – 7780 |
32 | J142232.38003043.9 | 2017-05-29 | 2600 | 18 | 18.5 | 5450 – 8450 |
2017-07-20 | 2500 | 336 | 17.75 | 5170 – 8190 | ||
33 | J142953.20+041551.6 | 2022-05-02 | 2300 | 76 | 19.625 | 5870 – 8840 |
2022-05-03 | 2300 | 166 | 19.625 | 5870 – 8840 | ||
34 | J161728.95+061604.3 | 2017-04-29 | 2300 | 45 | 20.75 | 6280 – 9250 |
2017-05-19 | 2137 | 330 | 20.75 | 6280 – 9250 | ||
35 | J215059.09+022557.2 | 2023-08-12 | 2500 | 175 | 18.125 | 5300 – 8300 |
36 | J223244.93+011058.1 | 2023-10-15 | 2500 | 130 | 17.0 | 4900 – 7900 |
37 | J224405.09021547.9 | 2023-06-23 | 2500 | 165 | 21.5 | 6560 – 9510 |
38 | J225028.85+031146.4 | 2023-07-17 | 2480 | 5 | 19.625 | 5870 – 8840 |
39 | J234257.50+010121.3 | 2023-07-22 | 2300 | 135 | 18.875 | 5590 – 8590 |
40 | J235639.31040614.4 | 2017-06-06 | 2500 | 45 | 20.375 | 6150 – 9120 |
2017-08-19 | 2350 | 315 | 20.375 | 6150 – 9120 | ||
2020-09-08 | 2560 | 90 | 20.0 | 6005 – 8985 |
The details of the observations of the GOTOQs systems are given in Table 2. The SALT spectroscopic observations were performed using the Robert Stobie Spectrograph (RSS) in long-slit mode from April 2017 to December 2023 (Program IDs: 2017-1-SCI-011, 2021-2-SCI-012, 2022-1-SCI-016, 2022-2-SCI-015, 2023-1-SCI-014, 2023-2-SCI-004). For our observations, we use the PG0900 grating along with a long slit-width of width and the grating angles are so chosen that the expected nebular lines from the GOTOQ host galaxies fall within the wavelength coverage of the spectrograph avoiding the CCD gaps.
The slit configurations used are shown in Figure 2 on top of the Dark Energy Camera Legacy Survey (DECaLS) r-band images of the quasar fields. SALT spectra were obtained for two or more PAs for 19 cases. In the remaining 21 cases we got spectra for one PA only. In ten of these sources the slit was aligned along the extensions seen in the images due to the presence of the foreground galaxies. For 11 sources (where there is no asymmetry clearly visible in the quasar images) the spectroscopic observations were not completed along the second PA.
The raw CCD frames obtained from the SALT observations were initially subjected to the SALT data reduction and analysis pipeline (Crawford et al., 2010). We then used the standard IRAF (Tody, 1986) routines to obtain the wavelength-calibrated 2D spectra and the flux-calibrated 1D spectra of the quasars. In summary, each science frame was first flat-field corrected, cosmic ray zapped, and then wavelength calibrated against a standard lamp. Then we applied the extinction correction due to the Earth’s atmosphere, and then the 1D spectra of the GOTOQs were extracted. The long-slit spectra were all individual flux calibrated against standard stars observed with the same settings as the GOTOQs. Once the flux calibration is done, we apply the air to vacuum wavelength transformation and also correct for the heliocentric velocity.
Our SALT observations reconfirm the presence of nebular [O ii] emission from foreground galaxies for 39 out of 40 systems observed. In the case of J1314+0330, we do not detect [O ii] emission. There are three SDSS spectra, obtained during three widely separated epochs, available for this source. The [O ii] emission is seen in only one spectrum that too in the one obtained with the narrower fiber. Therefore, we believe the detection of [O ii] emission in this case is false.
4 Results
4.1 Measuring the impact parameters
As mentioned earlier, the primary goal of this study is to measure the impact parameters of the GOTOQs using long-slit spectroscopy. One is looking for the offset of [O ii] emission with respect to the quasar trace at different position angles (PAs). We use these off-sets measured at different PAs to find the centroid of the [O ii] emission and hence the galaxy impact parameter (denoted as . Typically, one requires three PAs to have much better centroiding of the galaxy. However, due to observing time constraints, we could not get this for all the objects in our sample. However, we address this shortcoming by measuring the impact parameter (i.e., denoted by ) using the residual image after subtracting the quasar contribution whenever possible.
In the following, we illustrate the measurement for the GOTOQ towards J1209+0227. This is similar to what has been described in Noterdaeme et al. (2010). We begin with the reduced wavelength calibrated and background subtracted 2D SALT spectrum for the slit position angle of 0∘. We first model the quasar trace by a Gaussian in the spatial direction, whose amplitude is allowed to vary smoothly (a third-order polynomial) along the wavelength axis. The centroid of the Gaussian is allowed to vary linearly with the wavelength to account for the possible misalignment between the quasar trace and the pixels of the CCD. Then we subtract this model from the original 2D SALT spectrum to obtain the [O ii] emission without the quasar contamination as much as possible. We identify the centroid of this emission to measure the offset of the [O ii] emission relative to the background quasar trace. This process is illustrated in the top and middle left panels ( panel [A1] and [A2]) of Figure 3. The solid blue line corresponds to the quasar trace, while the red star marks the centroid of the [O ii] emission. The centroid is found to lie 0.4 pixels above the central quasar trace. We perform the same exercise for the spectra obtained at a position angle of 90∘, illustrated in the top and middle right panels ( panel [B1] and [B2]) of Figure 3, and the centroid is located 2.2 pixels above the central quasar trace. Once the centroids of the [O ii] emission are known along two different slit orientations, it is straightforward to measure the impact parameter of the GOTOQ. Since the background quasar and foreground galaxies are at very small angular separation, we use the plane geometry to find the impact parameter. Figure 3 illustrates this procedure in the bottom panels ( panel [C1] and [C2]). For this case we get kpc. For each GOTOQ, we also looked at the residual r-band image obtained after subtracting the Quasar PSF 111Detailed procedure for obtaining the PSF is given in https://psfex.readthedocs.io/en/latest/Working.html (as shown in the panel [C2] of Figure 3 for J1209+0227). We obtain kpc. The residual image also demonstrates the need for the spectra at the third PA to get the accurate impact parameters. We use the for statistical analysis in such cases. We similarly calculate the impact parameters for the rest of the GOTOQs in our sample. Note that when the residual flux is present after subtracting the quasar model, we attribute this residual flux to the foreground galaxy.
No. | System | ||
(kpc) | (kpc) | ||
(1) | (2) | (3) | (4) |
1 | J000123.61032317.9 | – | – |
2 | J005408.46094638.2 | 3.3 | 4.8 |
3 | J005436.35+005055.6 | – | 5.2 |
4 | J012916.14003211.4 | 13.6 | 13.5 |
5 | J013132.92040137.5 | – | 10.3 |
6 | J020317.22010122.8 | 7.3 | – |
7 | J023529.82+005916.3 | 9.3 | 6.3 |
8 | J023608.18+005604.0 | 7.5 | 6.3 |
9 | J023618.98000529.1 | – | 8.6 |
10 | J030730.61074555.6 | 1.2 | 7.4 |
11 | J080751.03+094554.7 | 15.2 | 16.3 |
12 | J080808.67+064108.7 | 8.6 | 9.2 |
13 | J084915.67+065013.0 | 10.5 | 9.5 |
14 | J085051.97+083026.6 | – | – |
15 | J085830.53+055606.0 | 7.0 | 7.7 |
16 | J090830.32+042436.3 | – | – |
17 | J091709.35+033734.0 | – | – |
18 | J092915.02012621.0 | 14.0 | 13.4 |
19 | J094222.51+001641.9 | – | – |
20 | J101446.63+083500.9 | – | 6.4 |
21 | J102518.88001755.6 | 10.0 | 11.2 |
22 | J111755.71+075932.2 | – | 8.4 |
23 | J120908.53+022734.3 | 3.8 | 5.9 |
24 | J121242.26+083410.9 | – | – |
25 | J121838.10+030900.1 | 9.3 | 11.4 |
26 | J123514.36+030416.7 | – | – |
27 | J131918.88030621.2 | – | 7.8 |
28 | J135445.72+075637.6 | 15.3 | 14.2 |
29 | J135634.16+060133.6 | 10.1 | 11.2 |
30 | J140446.28+062340.2 | – | 3.6 |
31 | J142232.38003043.9 | 8.1 | 9.0 |
32 | J142953.20+041551.6 | 13.9 | – |
33 | J161728.95+061604.3 | 3.9 | 9.8 |
34 | J215059.09+022557.2 | – | 5.7 |
35 | J223244.93+011058.1 | – | 4.4 |
36 | J224405.09021547.9 | 16.3 | 16.3 |
37 | J225028.85+031146.4 | – | – |
38 | J234257.50+010121.3 | 8.9 | 9.5 |
39 | J235639.31040614.4 | 6.2 | – |
The impact parameters obtained for our sample are listed in the third and fourth columns of Table 3. We measure in the case of 22 GOTOQs. Based on residual broadband images and the photometric extensions, we measure for 28 GOTOQs. For 19 GOTOQs we have measured both and . For 8 GOTOQs we could not measure either or . Spectra of six of these eight GOTOQs were obtained along one PA only. We believe the impact parameters are small in these cases as nebular emission lines are aligned with the quasar trace. For the two cases, J00010323 and J13190306, we could not constrain the impact parameters of the foreground galaxies even with spectra along two different PAs. For J00010323, although we detect [O ii] emission from the foreground galaxy along PA=0 aligning with the quasar trace, we do not detect any emission from the galaxy along PA=150, implying the foreground galaxy to be present along the PA of 60. For J13190306, the SNR of the quasar trace at the 45-degree position angle is too poor to determine the centroid of the [O ii] emission. We could measure the impact parameter for 31 GOTOQs using spectroscopic or photometric methods. In the case of J1235+0304, as the QSO image is blended with two nearby sources, we used only one slit position angle for our spectroscopic observations covering the quasar and those two sources. We detect a broad [O ii] emission exactly on top of the quasar. Our analysis confirms that the two blended objects seen in the image are not at the redshift of the GOTOQs and, hence, are not the host galaxies of the Mg ii absorption. No residual is detected when we model the image with three objects. In this case, we need spectra at another position angle to measure the impact parameter. Therefore, we do not include this object for further discussion. In three cases (GOTOQs towards J02030101, J23560406, and J1429+0415), we have measurements only based on long-slit spectroscopy as these systems have neither any photometric extension nor any residual left upon subtracting the quasar PSF. In the case of J23560406, the impact parameter measurement based on spectroscopy is secure as we use spectra obtained along three PAs. Among the systems with measured, for 12 cases, we have impact parameter measurements based on residual images rather than photometric extensions. However, in eight cases, we have spectroscopic confirmations of these residuals to be associated with foreground galaxies. For the remaining four cases, we have impact parameter measurements solely based on residual images.
For 19 GOTOQs, we have impact measurements based on both photometry and spectroscopy. We compare the two measurements in Figure 4. Apart from two measurements (i.e., for J03070745 and J1617+0616 shown in red), other measurements are consistent within kpc. In both these cases, the residual image is predominantly present in the regions not covered by our observations. We use the for all our analysis for these objects.
As expected, the measured impact parameters typically vary between 3 to 16 kpc with a median value of 8.6 kpc. For the median redshift (i.e., ) of or sample 1" corresponds to 7.2 kpc. From Table 1, we can find that 19 GOTOQs in our sample are only detected from the BOSS spectra where the fiber used has a diameter of 2". In only five cases, the measured impact parameters are less than 7.2 kpc. Therefore, nearly all our measurements of [O ii] luminosity in the BOSS spectra are affected by fiber losses. On the other hand, 10 out of 17 GOTOQs detected only in the SDSS spectra (with a fiber diameter of 3") have an impact parameter of less than 11 kpc. This implies that fiber loss effects may be less severe for these GOTOQs. There are 21 measurements with impact parameters less than ten kpc. This is by far the largest sample of Mg ii absorbers with spectroscopically confirmed host galaxies at such impact parameters. In what follows, we study the correlation between different absorption line parameters and impact parameters for this sample.
4.2 Rest equivalent widths versus the impact parameters
In this section, we study the relationship between the rest equivalent width (REW) of different species and impact parameters. For this we consider Mg ii 2796, Mg i 2852, Fe ii 2600, Ca ii and Mn ii . As expected, we detect strong Fe ii and Mg i absorption with at least 2 significance in almost all cases. We detect Ca ii 3935 and Mn ii 2576 absorption towards 16 and 17 GOTOQs in our observed sample, respectively. In Figure 5, we plot the rest equivalent width of all the species as a function of the impact parameter. The figure also shows the impact parameter corresponding to 1" and 1.5" using long- and short-dashed vertical lines at the median redshift. None of the species show any significant (anti-)correlation between the REW and the impact parameter.
4.2.1 Ca ii and Mg ii absorption:
Earlier studies (Wild et al., 2006; Sardane et al., 2015; Fang et al., 2023) suggested the existence of two populations of Ca ii absorbers defined based on the rest equivalent width of Ca ii (i.e., above (strong) and below (weak) 0.7Å). The strong Ca ii absorbers tend to produce stronger reddening signatures than weak Ca ii absorbers leading to the proposal that while strong Ca ii absorbers probe the inner dish+halo regions of galaxies, the weak Ca ii absorbers typically probe the halo (or CGM) of galaxies. One expects a possible anti-correlation between and the impact parameter in such cases. In Guha & Srianand (2023), using composite spectra, we showed the REW of Ca ii is higher for GOTOQs where the host galaxy is not well resolved compared to that of the resolved cases (i.e., their extended sample).
In our case, 9 out of 16 detections have Ca ii equivalent width Å. Thus, it belongs to the strong Ca ii absorbers population. This can be interpreted as the covering factor of strong Ca ii absorbers in the GOTOQs to be about 23%. In Figure 5, we do not find any correlation between REW of Ca ii and . A good fraction of upper limits occurs at smaller impact parameters. Typically, they are around 0.6Å due to poor SNR of the SDSS spectrum. This is consistent with our finding that of the USMgii and DLAs do not correlate with the impact parameter. Lack of correlation between the rest equivalent width of Ca ii and D was also reported by Straka et al. (2015) for their low-z GOTOQs. As in the case of Mg ii, one has to consider a larger impact parameter range to probe the dependence of Ca ii equivalent width and covering factor on the impact parameter.
Note that about 25% of our systems have consistent with them being defined as USMgii . In the detailed analysis presented in Guha et al. (2022) and Guha et al. (2024), we have shown that USMgii absorbers do not show any anti-correlation between and D. One of the possible reasons is that a good fraction of USMgii may originate from more than one galaxy. In the present case, it is most likely that the observed Mg ii absorption is associated with a single galaxy. From Figure 5, we do not find any correlation between and D. It is also evident from the figure that the USMgii systems and the remaining ones are distributed uniformly across the impact parameter probed.
In Guha & Srianand (2023), we have shown that including GOTOQs in the vs is crucial to distinguish between a simple power-law fit and a log-linear fit. However, in that work, the impact parameters for the GOTOQs were obtained, assuming that the galaxy identified in the DECaLS photometry is the real host. Here, we repeat the exercise considering only the GOTOQs studied here, focusing mainly on Mg ii absorbers at .
To compare the GOTOQs studied in our sample with other Mg ii absorption systems, we plot the same from the studies by Guha et al. (2022), Guha et al. (2024), Dutta et al. (2020); Huang et al. (2021); Nielsen et al. (2013). This is presented in Figure 6. The solid black line, which corresponds to the maximally likely log-linear fit to the combined observed data, is represented by the equation with an intrinsic scatter of and the gray shaded region corresponds to the error associated with the fit. The fitting procedure, including the upper limits, is described in detail in the appendix of Guha et al. (2022) and is similar to what is described in Chen et al. (2010); Dutta et al. (2020). The fit values are consistent with those in Guha et al. (2024).
4.3 Line of sight reddening towards GOTOQ
Assuming the Mg ii absorbing gas follows the dust extinction similar to that of the small magellanic clouds (SMC) (Gordon et al., 2003), we calculate the color excess of the background quasar spectra using a standard quasar spectral template (Selsing et al., 2016). We first redshift the quasar template to the redshift of the background quasar and then apply the SMC dust extinction law at the Mg ii absorption redshift to match it with the shape of the quasar spectrum in hand by varying the dust extinction coefficient, . The color excess, for our sample ranges from to 0.422, with a median value of 0.058. The measured values of are provided in the last column of Table 1. However, it is important to note that the highest line of sight reddening is observed for the system J1314+0330. As previously mentioned, this system is incorrectly identified as a GOTOQ. Therefore, like before, we do not include it in the analysis presented here. Given the median of 2.21 Å, the expected color excess towards this GOTOQ sub-sample is (Budzynski & Hewett, 2011). Therefore, the line of sight reddening towards the GOTOQ is significantly higher than the typical Mg ii absorption systems of similar . Wild & Hewett (2005) have measured the for a sample of Ca ii absorbers. Their values range from to 0.30 with a median value of 0.042.
In Figure 7, we compare the between the USMgii absorption systems at from Guha et al. (2022), Ca ii absorbers from Wild & Hewett (2005) and the GOTOQ sub-sample studied here. The blue and orange histograms correspond to the distribution of the GOTOQ and the USMgii absorption systems, respectively. A KS test between these two distributions yields a -value of 0.06, implying that the dust extinctions in these two samples follow similar distributions. This might be expected as one-fourth of the GOTOQs are consistent with being a USMgii system. Similarly, the -values obtained while comparing the E(BV) distribution for the GOTOQs and Ca ii systems is 0.24. This suggests that E(BV) distributions for these two samples are also statistically similar.
In Figure 8, we plot E(BV) vs. for GOTOQs discussed here, and for the Ca ii absorbers studied by Wild & Hewett (2005). We also compare these measurements of our Galaxy from Murga et al. (2015). Our measurements for GOTOQs roughly follow the region occupied by the Ca ii absorbers. However, we do not find any significant correlation when considering only our data points. It is evident from Figure 8 that for a given E(BV), the extra-galactic measurements have larger Ca ii rest equivalent width compared to the local measurements by Murga et al. (2015). While interpreting as an indicator of gas content is debatable, this could imply the dust-to-gas ratio in these absorbers is smaller than seen in our Galaxy. There are several indications for this whenever and (H i) measurements are available at high-z (see for example, discussions in Ledoux et al., 2015; Dutta et al., 2020). Lastly, we investigated the potential correlation between E(BV) and D in our sample and found no statistically significant correlation. The lack of correlation between E(BV) and impact parameter for the low-z GOTOQs was also reported by Straka et al. (2015).
4.4 Estimation of galaxy parameters
Galaxy | (Å) | |
(1) | (2) | (3) |
J01290032 | 3.500.26 | |
J01310401 | 1.700.12 | |
J0235+0059 | 0.930.43 | |
J0807+0945 | ||
J0808+0641 | ||
J0858+0556 | 2.901.03 | |
J09290126 | ||
J1218+0309 | 2.800.39 | |
J1354+0756 | 2.150.31 | |
J1356+0601 | 2.520.19 | |
J22440215 | 3.770.20 | |
J2342+0101 | 2.350.27 |
There is a correlation found between stellar mass and impact parameter for the general population of Mg ii absorbers in Guha et al. (2022). It has also been found that for a given impact parameter, the estimated for the USMgii absorbers tend to be higher compared to the Mg ii absorbers with low . These correlations imply that (i) at a given impact parameter, the absorbers with large tend to have large and (ii) for a given galaxies at higher impact parameters tend to have higher values of . Therefore, if GOTOQs follow this trend, we expect the for the host galaxies to be typically less than what has been found for USMgii host galaxies at high impact parameters. Indeed, Guha & Srianand (2023) have measured over the redshift range . Here, we check whether such measurements for individual cases in the case of GOTOQs follow the above-mentioned trend.
Accurate photometric measurements are required in several bands to measure the stellar masses () of GOTOQs using SED fitting techniques. This is possible only for 12 of the 39 objects with clear photometric extensions in our sample. For these twelve systems, the angular separations between the quasars and the foregrounds are large, so reliable fluxes in different photometric bands can be measured by subtracting the quasar PSF. We computed the foreground galaxy fluxes in the grz bands of the DECaLS photometry. Then, we fitted the spectral energy distributions (SED) using the publicly available code BAGPIPES (Carnall et al., 2018). During the fit, we assumed a delayed star formation history (Carnall et al., 2019) and kept the redshift constant as obtained from the [O ii] emission line. Obtained stellar masses of the GOTOQ with photometric counterparts are given in column 3 of Table 4. In Figure 9, we show plot vs. D for these objects and also show the best fit obtained for the host galaxies of USMgii and that obtained for the MAGIICAT galaxies. However, it is important to note that the correlation observed for the GOTOQs may be influenced by the mass-size relation of galaxies (Ward et al., 2024). Since more massive galaxies tend to be larger, massive GOTOQs are also larger in size and, consequently, may have larger impact parameters. It will be good to measure the host galaxy properties for more GOTOQs using HST photometry or AO-assisted ground-based imaging. This will provide more insights into how stellar mass controls the equivalent width distributions at low impact parameters.
5 Summary
In this work, we present long-slit spectroscopic observations of 40 GOTOQs having [O ii] luminosity above over the redshift range . The Mg ii rest equivalent width of our sample spans Å with a median value of 2.21Å.
We are successful in measuring the impact parameters for 31 GOTOQs. The impact parameter ranges from 3.6 to 16.3 kpc with a median impact parameter of 8.6 kpc. In 19 cases, we have simultaneous impact parameter measurements using our spectroscopic observations with SALT and photometric observations from DECaLS. Apart from the two cases, they agree very well between them within an uncertainty of 1.5 kpc. This reassures that the residual fluxes measured around quasars with correct photometric redshifts are the host galaxies of GOTOQs. The measured impact parameters also confirm that the [O ii] luminosity measured using SDSS fiber spectra is severely affected by fiber losses, particularly the BOSS spectra.
We do not find any correlation or anti-correlation between and D. This aligns with our earlier finding that these two quantities are uncorrelated at small impact parameters. Based on all the available spectroscopically confirmed host galaxy measurements, we provide the best-fitted log-linear relationship between these two quantities.
We report the detection of Ca ii 3935 and Mn ii 2576 towards 16 and 17 GOTOQs, respectively. For Å (i.e., strong Ca ii absorbers by usual definition), we find the covering factor to be 0.23 around GOTOQs. We do not find any correlation between and D. This is consistent with the finding of Straka et al. (2015) for low-z GOTOQs. It is usually conjectured that weak Ca ii absorbers originate mainly from the halo gas, while strong Ca ii absorbers originate from the disk and/or inner halo gas. In such a case, we expect an anti-correlation between and D. The Absence of this correlation could be related to the lack of measurements from absorbers at higher D having low . We also explore the correlation between REW of Mn ii 2576, Fe ii 2600 and Mg i 2852. We do not find any statistically significant correlation or anti-correlation with D for REW of these ions.
The line of sight reddening (quantified using E(BV)) obtained using the SMC extinction curve for GOTOQs is statistically similar for high-z Mg ii systems. They are also found to be similar to what has been measured for Ca ii selected absorbers studied by Wild & Hewett (2005). In the E(BV) vs plane, GOTOQs occupy the same region as Ca ii absorbers of Wild & Hewett (2005). However, for a given E(BV), both the sample show larger than what has been found in our Galaxy by Murga et al. (2015). This can be attributed to the smaller dust-to-gas ratio in high- absorbers. We do not find any correlation between E(BV) and impact parameters. This is also consistent with the result of Straka et al. (2015) for low-z GOTOQs.
Measurement of is possible for 12 GOTOQs in our sample. As in the case of the general population of Mg ii absorbers, these measurements are consistent with (i) at a given impact parameter, the absorbers with large tend to have large and (ii) for a given galaxies at higher impact parameters tend to have higher values of . Establishing the relationship between these quantities using a large sample is essential for understanding how stellar mass controls the equivalent width distributions at low impact parameters.
Data Availability
Data used in this work are obtained using SALT. Raw data will become available for public use 1.5 years after the observing date at https://ssda.saao.ac.za/.
Acknowledgement
This project makes use of the following softwares : NumPy (Harris et al., 2020), SciPy (Virtanen et al., 2020), Matplotlib (Hunter, 2007), AstroPy (Astropy Collaboration et al., 2013, 2018), and Ultranest (Buchner, 2021).
All the new observations reported in this paper were obtained with the Southern African Large Telescope (SALT).
This paper makes use of SDSS observational data. Funding for the Sloan Digital Sky Survey IV has been provided by the Alfred P. Sloan Foundation, the U.S. Department of Energy Office of Science, and the Participating Institutions. SDSS-IV acknowledges support and resources from the Center for High-Performance Computing at the University of Utah. The SDSS website is www.sdss.org. SDSS-IV is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration, including the Brazilian Participation Group, the Carnegie Institution for Science, Carnegie Mellon University, Center for Astrophysics | Harvard & Smithsonian, the Chilean Participation Group, the French Participation Group, Instituto de Astrofísica de Canarias, The Johns Hopkins University, Kavli Institute for the Physics and Mathematics of the Universe (IPMU) / University of Tokyo, the Korean Participation Group, Lawrence Berkeley National Laboratory, Leibniz Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Astrophysik (MPA Garching), Max-Planck-Institut für Extraterrestrische Physik (MPE), National Astronomical Observatories of China, New Mexico State University, New York University, University of Notre Dame, Observatário Nacional / MCTI, The Ohio State University, Pennsylvania State University, Shanghai Astronomical Observatory, United Kingdom Participation Group, niversidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Oxford, University of Portsmouth, University of Utah, University of Virginia, University of Washington, University of Wisconsin, Vanderbilt University, and Yale University.
The DESI Legacy Imaging Surveys consist of three individual and complementary projects: the Dark Energy Camera Legacy Survey (DECaLS), the Beijing-Arizona Sky Survey (BASS), and the Mayall z-band Legacy Survey (MzLS). DECaLS, BASS, and MzLS together include data obtained, respectively, at the Blanco telescope, Cerro Tololo Inter-American Observatory, NSF’s NOIRLab; the Bok telescope, Steward Observatory, University of Arizona; and the Mayall telescope, Kitt Peak National Observatory, NOIRLab. NOIRLab is operated by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation. Pipeline processing and analyses of the data were supported by NOIRLab and the Lawrence Berkeley National Laboratory (LBNL). Legacy Surveys also uses data products from the Near-Earth Object Wide-field Infrared Survey Explorer (NEOWISE), a project of the Jet Propulsion Laboratory/California Institute of Technology, funded by the National Aeronautics and Space Administration. Legacy Surveys was supported by: the Director, Office of Science, Office of High Energy Physics of the U.S. Department of Energy; the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility; the U.S. National Science Foundation, Division of Astronomical Sciences; the National Astronomical Observatories of China, the Chinese Academy of Sciences and the Chinese National Natural Science Foundation. LBNL is managed by the Regents of the University of California under contract to the U.S. Department of Energy. The complete acknowledgments can be found at https://www.legacysurvey.org/acknowledgment/. The Photometric Redshifts for the Legacy Surveys (PRLS) catalog used in this paper was produced thanks to funding from the U.S. Department of Energy Office of Science, Office of High Energy Physics via grant DE-SC0007914.